QUEERS READ THIS A leaflet distributed at pride march in NY

Published anonymously by Queers

June, 1990

How can I tell you. How can I convince you, brother, sister that your life is in danger: That everyday you wake up alive, relatively happy, and a functioning human being, you are committing a rebellious act. You as an alive and functioning queer are a revolutionary. There is nothing on this planet that validates, protects or encourages your existence. It is a miracle you are standing here reading these words. You should by all rights be dead.

Don’t be fooled, straight people own the world and the only reason you have been spared is you’re smart, lucky or a fighter. Straight people have a privilege that allows them to do whatever they please and fuck without fear. But not only do they live a life free of fear; they flaunt their freedom in my face. Their images are on my TV, in the magazine I bought, in the restaurant I want to eat in, and on the street where I live. I want there to be a moratorium on straight marriage, on babies, on public displays of affection among the opposite sex and media images that promote heterosexuality. Until I can enjoy the same freedom of movement and sexuality, as straights, their privilege must stop and it must be given over to me and my queer sisters and brothers.

Straight people will not do this voluntarily and so they must be forced into it. Straights must be frightened into it. Terrorized into it. Fear is the most powerful motivation. No one will give us what we deserve. Rights are not given they are taken, by force if necessary.

It is easier to fight when you know who your enemy is. Straight people are your enemy. They are your enemy when they don’t acknowledge your invisibility and continue to live in and contribute to a culture that kills you.

Every day one of us is taken by the enemy. Whether it’s an AIDS death due to homophobic government inaction or a lesbian bashing in an all-night diner (in a supposedly lesbian neighborhood), [Ed. note: a reference to an incident which occurred in Park Slope, Brooklyn.] we are being systematically picked off and we will continue to be wiped out unless we realize that if they take one of us they must take all of us.

AN ARMY OF LOVERS CANNOT LOSE

Being queer is not about a right to privacy; it is about the freedom to be public, to just be who we are. It means everyday fighting oppression; homophobia, racism, misogyny, the bigotry of religious hypocrites and our own self-hatred. (We have been carefully taught to hate ourselves.) And now of course it means fighting a virus as well, and all those homo-haters who are using AIDS to wipe us off the face of the earth.

Being queer means leading a different sort of life. It’s not about the mainstream, profit-margins, patriotism, patriarchy or being assimilated. It’s not about executive directors, privilege and elitism. It’s about being on the margins, defining ourselves; it’s about gender-fuck and secrets, what’s beneath the belt and deep inside the heart; it’s about the night. Being queer is “grass roots” because we know that everyone of us, every body, every cunt, every heart and ass and dick is a world of pleasure waiting to be explored. Everyone of us is a world of infinite possibility. We are an army because we have to be. We are an army because we are so powerful. (We have so much to fight for; we are the most precious of endangered species.) And we are an army of lovers because it is we who know what love is. Desire and lust, too. We invented them. We come out of the closet, face the rejection of society, face firing squads, just to love each other! Every time we fuck, we win.

We must fight for ourselves (no one else is going to do it) and if in that process we bring greater freedom to the world at large then great. (We’ve given so much to that world: democracy, all the arts, the concepts of love, philosophy and the soul, to name just a few gifts from our ancient Greek Dykes, Fags.) Let’s make every space a Lesbian and Gay space. Every street a part of our sexual geography. A city of yearning and then total satisfaction. A city and a country where we can be safe and free and more. We must look at our lives and see what’s best in them, see what is queer and what is straight and let that straight chaff fall away! Remember there is so, so little time. And I want to be a lover of each and every one of you. Next year, we march naked.

ANGER

“The strong sisters told the brothers that there were two important things to remember about the coming revolutions, the first is that we will get our asses kicked. The second, is that we will win.”

I’m angry. I’m angry for being condemned to death by strangers saying, “You deserve to die” and “AIDS is the cure.” Fury erupts when a Republican woman wearing thousands of dollars of garments and jewelry minces by the police lines shaking her head, chuckling and wagging her finger at us like we are recalcitrant children making absurd demands and throwing temper tantrum when they aren’t met. Angry while Joseph agonizes over $8,000 a over for AZT which might keep him alive a little longer and which makes him sicker than the disease he is diagnosed with. Angry as I listen to a man tell me that after changing his will five times he’s running out of people to leave things to. All of his best friends are dead. Angry when stand in a sea of quilt panels, or go to a candlelight march or attend yet another memorial service. I will not march silently with a fucking candle and I want to take that goddamned quilt and wrap myself in it and furiously rend it and my hair and curse every god religion ever created. I refuse to accept a creation that cuts people down in the third decade of their life. It is cruel and vile and meaningless and everything I have in me rails against the absurdity and I raise my face to the clouds and a ragged laugh that sounds more demonic than joyous erupts from my throat and tears stream down my face and if this disease doesn’t kill me, I may just die of frustration. My feet pound the streets and Peter’s hands are chained to a pharmaceutical company’s reception desk while the receptionist looks on in horror and Eric’s body lies rotting in a Brooklyn cemetery and I’ll never hear his flute resounding off the walls of the meeting house again. And I see the old people in Tompkins Square Park huddled in their long wool coats in June to keep out the cold they perceive is there and to cling to whatever little life has left to offer them. I’m reminded of the people who strip and stand before a mirror each night before they go to bed and search their bodies for any mark that might not have been there yesterday. A mark that this scourge has visited them. And I’m angry when the newspapers call us “victims” and sound alarms that “it” might soon spread to the “general population.” And I want to scream “Who the fuck am I?” And I want to scream at New York Hospital with its yellow plastic bags marked “isolation linen,” “ropa infecciosa” and its orderlies in latex gloves and surgical masks skirting the bed as if its occupant will suddenly leap out and douse them with blood and semen giving them too the plague. And I’m angry at straight people who sit smugly wrapped in their self-protective coat of monogamy and heterosexuality confident that this disease has nothing to do with them because “it” only happens to “them.” And the teenage boys who upon spotting my Silence=Death button begin chanting “Faggot’s gonna die” and I wonder, who taught them this? Enveloped in fury and fear, I remain silent while my button mocks me every step of the way. And the anger I fell when a television program on the quilt gives profiles of the dead and the list begins with a baby, a teenage girl who got a blood transfusion, an elderly baptist minister and his wife and when they finally show a gay man, he’s described as someone who knowingly infected teenage male prostitutes with the virus. What else can you expect from a faggot? I’m angry.

QUEER ARTISTS

Since time began, the world has been inspired by the work of queer artists. In exchange, there has been suffering, there has been pain, there has been violence. Throughout history, society has struck a bargain with its queer citizens: they may pursue creative careers, if they do it discreetly. Through the arts queers are productive, lucrative, entertaining and even uplifting. These are the clear-cut and useful by-products of what is otherwise considered antisocial behavior. In cultured circles, queers may quietly coexist with an otherwise disapproving power elite.

At the forefront of the most recent campaign to bash queer artists is Jesse Helms, arbiter of all that is decent, moral, christian and amerikan. For Helms, queer art is quite simply a threat to the world. In his imaginings, heterosexual culture is too fragile to bear up to the admission of human or sexual diversity. Quite simply, the structure of power in the Judeo-Christian world has made procreation its cornerstone. Families having children assures consumers for the nation’s products and a work force to produce them, as well as a built-in family system to care for its ill, reducing the expense of public healthcare systems. ALL NON-PROCREATIVE BEHAVIOR IS CONSIDERED A THREAT, from homosexuality to birth control to abortion as an option. It is not enough, according to the religious right, to consistently advertise procreation and heterosexuality… it is also necessary to destroy any alternatives. It is not art Helms is after… IT IS OUR LIVES! Art is the last safe place for lesbians and gay men to thrive. Helms knows this, and has developed a program to purge queers from the one arena they have been permitted to contribute to our shared culture.

Helms is advocating a world free from diversity or dissent. It is easy to imagine why that might feel more comfortable to those in charge of such a world. It is also easy to envision an amerikan landscape flattened by such power. Helms should just ask for what he is hinting at: State sponsored art, art of totalitarianism, art that speaks only in christian terms, art which supports the goals of those in power, art that matches the sofas in the Oval Office. Ask for what you want, Jesse, so that men and women of conscience can mobilize against it, as we do against the human rights violations of other countries, and fight to free our own country’s dissidents.

IF

YOU’RE QUEER,

Queers are under siege.

Queers are being attacked on all fronts and I’m afraid it’s ok with us.

In 1969, there were 50 “Queer Bashings” in the month of May alone. Violent attacks, 3 720 men, women and children died of AIDS in the same month, caused by a more violent attack — government inaction, rooted in society’s growing homophobia. This is institutionalized violence, perhaps more dangerous to the existence of queers because the attackers are faceless. We allow these attacks by our own continued lack of action against them. AIDS has affected the straight world and now they’re blaming us for AIDS and using it as a way to justify their violence against us. They don’t want us anymore. They will beat us, rape us and kill us before they will continue to live with us.

What will it take for this not to be ok? Feel some rage. If rage doesn’t empower you, try fear. If that doesn’t work, try panic.

SHOUT IT!

Be proud. Do whatever you need to do to tear yourself away from your customary state of acceptance. Be free. Shout.

In 1969, Queers fought back. In 1990, Queers say ok. Next year, will we be here?

I HATE…

I hate Jesse Helms. I hate Jesse Helms so much I’d rejoice if he dropped down dead. If someone killed him I’d consider it his own fault.

I hate Ronald Reagan, too, because he mass-murdered my people for eight years. But to be honest, I hate him even more for eulogizing Ryan White without first admitting his guilt, without begging forgiveness for Ryan’s death and for the deaths of tens of thousands of other PWA’s — most of them queer. I hate him for making a mockery of our grief.

I hate the fucking Pope, and I hate John fucking Cardinal fucking O’Connor, and I hate the whole fucking Catholic Church. The same goes for the Military, and especially for Amerika’s Law Enforcement Officials — the cops — state sanctioned sadists who brutalize street transvestites, prostitutes and queer prisoners. I also hate the medical and mental health establishments, particularly the psychiatrist who conviced me not to have sex with men for three years until we (meaning he) could make me bisexual rather than queer. I also hate the education profession, for its share in driving thousands of queer teens to suicide every year. I hate the “respectable” art world; and the entertainment industry, and the mainstream media, especially The New York Times. In fact, I hate every sector of the straight establishment in this country — the worst of whom actively want all queers dead, the best of whom never stick their necks out to keep us alive.

I hate straight people who think they have anything intelligent to say about “outing.” I hate straight people who think stories about themselves are “universal” but stories about us are only about homosexuality. I hate straight recording artists who make their careers off of queer people, then attack us, then act hurt when we get angry and then deny having wronged us rather than apologize for it. I hate straight people who say, “I don’t see why you feel the need to wear those buttons and t-shirts. I don’t go around telling the whole world I’m straight.”

I hate that in twelve years of public education I was never taught about queer people. I hate that I grew up thinking I was the only queer in the world, and I hate even more that most queer kids still grow up the same way. I hate that I was tormented by other kids for being a faggot, but more that I was taught to feel ashamed for being the object of their cruelty, taught to feel it was my fault. I hate that the Supreme Court of this country says it’s okay to criminalize me because of how I make love. I hate that so many straight people are so concerned about my goddamned sex life. I hate that so many twisted straight people become parents, while I have to fight like hell to be allowed to be a father. I hate straights.

WHERE ARE YOU SISTERS?

I wear my pink triangle everywhere. I do not lower my voice in public when talking about lesbian love or sex. I always tell people I’m a lesbian. I don’t wait to be asked about my “boyfriend.” I don’t say it’s “no one’s business.”

I don’t do this for straight people. Most of them don’t know what the pink triangle even means. Most of them couldn’t care less that my girlfriend and I are totally in love or having a fight on the street. Most of them don’t notice us no matter what we do. I do what I do to reach other lesbians. I do what I do because I don’t want lesbians to assume I’m a straight girl. I am out all the time, everywhere, because I WANT TO REACH YOU. Maybe you’ll notice me, maybe we’ll start talking, maybe we’ll exchange numbers, maybe we’ll become friends. Maybe we won’t say a word but our eyes will meet and I will imagine you naked, sweating, openmouthed, your back arched as I am fucking you. And we’ll be happy to know we aren’t the only ones in the world. We’ll be happy because we found each other, without saying a word, maybe just for a moment.

But no.

You won’t wear a pink triangle on that linen lapel. You won’t meet my eyes if I flirt with you on the street. You avoid me on the job because I’m “too” out. You chastise me in bars because I’m “too political.” You ignore me in public because I bring “too much” attention to “my” lesbianism. But then you want me to be your lover, you want me to be your friend, you want me to love you, support, you, fight for “OUR”r ight to exist.

WHERE ARE YOU?

You talk, talk, talk about invisibility and then retreat to your homes to nest with your lovers or carouse in a bar with pals and stumble home in a cab or sit silently and politely by while your family, your boss, your neighbors, your public servants distort and disfigure us, deride us and punish us. Then home again and you feel like screaming. Then you pad your anger with a relationship or a career or a party with other dykes like you and still you wonder why we can’t find each other, why you feel lonely, angry, alienated.

GET UP, WAKE UP SISTERS!!

Your life is in your hands.

When I risk it all to be out, I risk it for both of us. When I risk it all and it works (which it often does if you would try it), I benefit and so do you. When it doesn’t work, I suffer and you do not.

But girl you can’t wait for other dykes to make the world safe for you. STOP waiting for a better more lesbian future! The revolution could be here if we started it.

Where are you sisters? I’m trying to find you, I’m trying to find you. How come I only see you on Gay Pride Day?

We’re OUT, Where the fuck are YOU?

WHEN ANYONE ASSAULTS YOU FOR BEING QUEER, IT IS QUEER BASHING. RIGHT?

A crowd of 50 people exit a gay bar as it closes. Across the street, some straight boys are shouting “Faggots” and throwing beer bottles at the gathering, which outnumbers them by 10 to 1. Three queers make a move to respond, getting no support from the group. Why did a group this size allow themselves to be sitting ducks?

Tompkins Square Park, Labor Day. At an annual outdoor concert/drag show, a group of gay men were harassed by teens carrying sticks. In the midst of thousands of gay men and lesbians, these straight boys beat two gay men to the ground, then stood around triumphantly laughing amongst themselves. The emcee was alerted and warned the crowd from the stage, “You girls be careful. When you dress up it drives the boys crazy,” as if it were a practical joke inspired by what the victims were wearing rather than a pointed attack on anyone and everyone at that event.

What would it have taken for that crowd to stand up to its attackers?

After James Zappalorti, an openly gay man, was murdered in cold blood on Staten Island this winter, a single demonstration was held in protest. Only one hundred people came. When Yuseuf Hawkins, a black youth, was shot to death for being on “white turf” in Bensonhurst, African Americans marched through that neighborhood in large numbers again and again. A black person was killed BECAUSE HE WAS BLACK, and people of color throughout the city recognized it and acted on it. The bullet that hit Hawkins was meant for a black man, ANY black man. Do most gays and lesbians think that the knife that punctured Zappalorti’s heart was meant only for him?

The straight world has us so convinced that we are helpless and deserving victims of the violence against us, that queers are immobilized when faced with a threat. BE OUTRAGED! These attacks must not be tolerated. DO SOMETHING. Recognize that any act of aggression against any member of our community is an attack on every member of the community. The more we allow homophobes to inflict violence, terror and fear on our lives, the more frequently and ferociously we will be the object of their hatred. Your body cannot be an open target for violence. Your body is worth protecting. You have a right to defend it. No matter what they tell you, your queerness must be defended and respected. You’d better learn that your life is immeasurably valuable, because unless you start believing that, it can easily be taken from you. If you know how to gently and efficiently immobilize your attacker, then by all means, do it. If you lack those skills, then think about gouging out his fucking eyes, slamming his nose back into his brain, slashing his throat with a broken bottle — do whatever you can, whatever you have to, to save your life!

reeuQ yhW

Queer!

Ah, do we really have to use that word? It’s trouble. Every gay person has his or her own take on it. For some it means strange and eccentric and kind of mysterious. That’s okay, we like that. But some gay girls and boys don’t. They think they’re more normal than strange. And for others “queer” conjures up those awful memories of adolescent suffering. Queer. It’s forcibly bittersweet and quaint at best — weakening and painful at worst. Couldn’t we just use “gay” instead? It’s a much brighter word and isn’t it synonymous with “happy?” When will you militants grow up and get over the novelty of being different?

WHY QUEER

Well, yes, “gay ” is great. It has its place. But when a lot of lesbians and gay men wake up in the morning we feel angry and disgusted, not gay. So we’ve chosen to call ourselves queer. Using “queer” is a way of reminding us how we are perceived by the rest of the world. It’s a way of telling ourselves we don’t have to be witty and charming people who keep our lives discreet and marginalized in the straight world. We use queer as gay men loving lesbians and lesbians loving being queer. Queer, unlike GAY, doesn’t mean MALE.

And when spoken to other gays and lesbians it’s a way of suggesting we close ranks, and forget (temporarily) our individual differences because we face a more insidious common enemy. Yeah, QUEER can be a rough word but it is also a sly and ironic weapon we can steal from the homophobe’s hands and use against him.

NO SEX POLICE

For anyone to say that coming out is not part of the revolution is missing the point. Positive sexual images and what they manifest saves lives because they affirm those lives and make it possible for people to attempt to live as self-loving instead of self-loathing. As the famous “Black is beautiful” slogan changed many lives, so does “Read my lips” affirm queerness in the face of hatred and invisibility as displayed in a recent governmental study of suicides that states at least one third of all teen suicides are Queer kids. This is further exemplified by the rise in HIV transmission among those under 21.

We are most hated as queers for our sexualness, that is, our physical contact with the same sex. Our sexuality and sexual expression are what makes us most susceptible to physical violence. Our difference, our otherness, our uniqueness can either paralyze us or politicize us. Hopefully, the majority of us will not let it kill us.

QUEER SPACE

Why in the world do we let heteros into queer clubs? Who gives a fuck if they like us because we “really know how to party?” WE HAVE TO IN ORDER TO BLOW OFF THE STEAM THEY MAKE US FEEL ALL THE TIME! They make out wherever they please, and take up too much room on the dance floor doing ostentatious couples dances. They wear their heterosexuality like a “Keep Out” sign, or like a deed of ownership.

Why the fuck do we tolerate them when they invade our space like it’s their right? Why do we let them shove heterosexuality — a weapon their world wields against us — right in our faces in the few public spots where we can be sexy with each other and not fear attack?

It’s time to stop letting the straight people make all the rules. Let’s start by posting this sign outside every queer club and bar:

RULES OF CONDUCT FOR STRAIGHT PEOPLE

1. Keep your display of affection (kissing, handholding, embracing) to a minimum. Your sexuality is unwanted and offensive to many here.

2. If you must slow dance, be as inconspicuous as possible.

3. Do not gawk or stare at lesbians or gay men, especially bull dykes or drag queens. We are not your entertainment.

4. If you cannot comfortably deal with someone of the same sex making a pass at you, get out.

5. Do not flaunt your heterosexuality. Be Discreet. Risk being mistaken for a lezzie or a homo.

6. If you feel these rules are unfair, go fight homophobia in straight clubs, or

7. Go Fuck Yourself.

I HATE STRAIGHTS

I have friends. Some of them are straight.

Year after year, I see my straight friends. I want to see them, to see how they are doing, to add newness to our long and complicated histories, to experience some continuity.

Year after year I continue to realize that the facts of my life are irrelevant to them and that I am only half listened to, that I am an appendage to the doings of a greater world, a world of power and privilege, of the laws of installation, a world of exclusion. “That’s not true,” argue my straight friends. There is the one certainty in the politics of power: those left out of it beg for inclusion, while the insiders claim that they already are. Men do it to women, whites do it to blacks, and everyone does it to queers.

The main dividing line, both conscious and unconscious, is procreation… and that magic word — Family. Frequently, the ones we are born into disown us when they find out who we really are, and to make matters worse, we are prevented from having our own. We are punished, insulted, cut off, and treated like seditionaries in terms of child rearing, both damned if we try and damned if we abstain. It’s as if the propagation of the species is such a fragile directive that without enforcing it as if it were an agenda, humankind would melt back into the primeval ooze.

I hate having to convice straight people that lesbians and gays live in a war zone, that we’re surrounded by bomb blasts only we seem to hear, that our bodies and souls are heaped high, dead from fright or bashed or raped, dying of grief or disease, stripped of our personhood.

I hate straight people who can’t listen to queer anger without saying “hey, all straight people aren’t like that. I’m straight too, you know,” as if their egos don’t get enough stroking or protection in this arrogant, heterosexist world. Why must we take care of them, in the midst of our just anger brought on by their fucked up society?! Why add the reassurance of “Of course, I don’t mean you. You don’t act that way.” Let them figure out for themselves whether they deserve to be included in our anger.

But of course that would mean listening to our anger, which they almost never do. They deflect it, by saying “I’m not like that” or “Now look who’s generalizing” or “You’ll catch more flies with honey… ” or “If you focus on the negative you just give out more power” or “you’re not the only one in the world who’s suffering.” They say “Don’t yell at me, I’m on your side” or “I think you’re overreacting” or “BOY, YOU’RE BITTER.”

They’ve taught us that good queers don’t get mad. They’ve taught us so well that we not only hide our anger from them, we hide it from each other. WE EVEN HIDE IT FROM OUR-SELVES. We hide it with substance abuse and suicide and overarhcieving in the hope of proving our worth. They bash us and stab us and shoot us and bomb us in ever increasing numbers and still we freak out when angry queers carry banners or signs that say BASH BACK. For the last decade they let us die in droves and still we thank President Bush for planting a fucking tree, applaud him for likening PWAs to car accident victims who refuse to wear seatbelts. LET YOURSELF BE ANGRY. Let yourself be angry that the price of our visibility is the constant threat of violence, anti-queer violence to which practically every segment of this society contributes. Let yourself feel angry that THERE IS NO PLACE IN THIS COUNTRY WHERE WE ARE SAFE, no place where we are not targeted for hatred and attack, the self-hatred, the suicide — of the closet. The next time some straight person comes down on you for being angry, tell them that until things change, you don’t need any more evidence that the world turns at your expense. You don’t need to see only hetero couple grocery shopping on your TV… You don’t want any more baby pictures shoved in your face until you can have or keep your own. No more weddings, showers, anniversaries, please, unless they are our own brothers and sisters celebrating. And tell them not to dismiss you by saying “You have rights,” “You have privileges,” “You’re overreacting,” or “You have a victim’s mentality.” Tell them “GO AWAY FROM ME, until YOU can change.” Go away and try on a world without the brave, strong queers that are its backbone, that are its guts and brains and souls. Go tell them go away until they have spent a month walking hand in hand in public with someone of the same sex. After they survive that, then you’ll hear what they have to say about queer anger. Otherwise, tell them to shut up and listen.